# Simulating Optical Coherent Nonlinear Response for High-Intensity Excitation

### v1

@@ Abstract
Calculation of the coherent nonlinear response of a system is essential to correctly interpret results from advanced techniques such as two-dimensional coherent spectroscopy (2DCS). Usually, even for the simplest systems, such calculations are either performed for low-intensity excitations where perturbative methods are valid and/or by assuming a simplified pulse envelope, such as a $\delta$-function in time. We present exact calculations using the phase-cycling method without making the aforementioned approximations. We introduce a generalized version of the phase-cycling method to isolate an arbitrary $N$-wave mixing signal. We then apply this method to model the saturation of the nonlinear signal from excitons in semiconductor quantum wells, which is consistent with 2DCS experiments. We also present simulation results that replicate previously-reported experiments with high-intensity excitation of semiconductor quantum dots. By accurately reproducing a variety of phenomena such as higher-order contributions, switching of coherent signal, and changes in photon-echo transients, we prove the efficacy of the phase-cycling method to calculate the coherent nonlinear signal for high-intensity excitation. This method would be particularly useful for systems with multiple, well-separated peaks and/or large inhomogeneity.
@@ I Introduction
Nonlinear coherent spectroscopy is a widely used technique for probing complex light-matter interactions at the quantum level Mukamel (1995); Cundiff (2008). A fundamental component of nonlinear spectroscopy is the wave-mixing process Shen (2003); Bokor et al. (1981), where a sequence of laser pulses interacts with the material to generate nonlinear signals based on the relevant nonlinear susceptibility of the material. Among these interactions, four-wave mixing (FWM) is one of the most common, arising from the third-order susceptibility $\chi^{(3)}$ Vaughan et al. (2007). Higher-order susceptibilities ($\chi^{(n)}$) lead to wave-mixing processes such as six-wave mixing (SWM) and beyond Zhang et al. (2013); Gibson et al. (1991); Trebino and Rahn (1987); Tominaga and Yoshihara (1996); Steffen and Duppen (1996); Lucht, Trebino, and Rahn (1992); Kang, Hernandez, and Zhu (2004), which have been used to probe the electronic Oliver and Fleming (2015) and vibrational states Bhattacharyya and Fleming (2019); Liu et al. (2019) of the system.
@@
Building on the wave mixing principle, multidimensional coherent spectroscopy (MDCS) extends this approach by measuring the wave-mixing signal field while varying two or more time delays between interacting excitation pulses Li et al. (2023); Oliver (2018); Gallagher Faeder and Jonas (1999); Cho (2008); Brixner et al. (2017); Hamm and Zanni (2011); Zhang et al. (2023, 2013); Zhang, Wells, and Tan (2012). By performing Fourier transforms along the axes of these time delays, MDCS disentangles the nonlinear interactions of the material across multiple frequency dimensions. This multidimensional approach offers several advantages: it measures spectrally-resolved homogeneous response, isolates individual resonances in congested spectrum, and reveals the dynamic evolution trajectories of states that would otherwise remain hidden in one-dimensional measurements. As a result, MDCS provides in-depth insights into phenomena such as energy transfer Thyrhaug et al. (2016); Bixner et al. (2012); Bressan et al. (2019), chemical reactions Nuernberger, Ruetzel, and Brixner (2015), and quantum coherence processes Butkus et al. (2017); Turner and Nelson (2010).
@@
The spectra obtained from MDCS are rich in information, revealing intricate features like coherent couplings Singh et al. (2014); Li et al. (2013a); Wen and Nelson (2013) and many-body interactions Shacklette and Cundiff (2003); Erementchouk, Leuenberger, and Sham (2007). Proper interpretation of these complex signals requires robust simulations. Conventionally, theoretical and numerical models for simulation rely on a perturbative framework, where the nonlinear response scales linearly with the electric field of each excitation pulse, for weak excitation. In this framework, Feynman diagrams (FDs) are often employed to map out specific quantum pathways Yang et al. (2007); Rose and Krich (2021); Tekavec, Lott, and Marcus (2007), while response functions integrated with electric fields are used to compute the nonlinear response Mukamel (1995); Gelin et al. (2013). However, the overall response is ultimately truncated to a specific order, as it assumes the perturbative expansion remains valid Gelin, Egorova, and Domcke (2009); Gelin et al. (2003); Süß et al. (2019). Consequently, MDCS experiments are generally conducted using low intensity laser pulses, ensuring that higher-order contributions remain below the noise level of measurements, keeping the analysis within the bounds of perturbation theory.
@@
On the other hand, high-intensity excitation can significantly influence coherent excitation and population transfer dynamics, as higher-order NWM processes become significant. MDCS experiments have demonstrated that intense laser field reveals various interesting phenomena across different systems, such as correlations among multiple particles Liang and Li (2021), coherent control schemes through strong-field coupling Suzuki et al. (2016); Wen and Nelson (2013); Cerullo et al. (1996); Schneider et al. (2011), coherence between highly excited states in quantum wellsTurner and Nelson (2010), and exciton-exciton annihilation (EEA) processes in excitonic systems Malý and Mančal (2018); Trinkunas et al. (2001). However, most theoretical descriptions of these phenomena rely on simplified approximations such as approximate electric-field envelopes of the excitation pulses and/or truncation of the signal up to a particular nonlinear order Anda and Cole (2021); Binz et al. (2020); Bayer et al. (2009); Chen et al. (2017), which may not be valid for all cases. Moreover, several studies have highlighted that finite pulse durations and overlaps can affect the MDCS spectra within the perturbative regime Smallwood, Autry, and Cundiff (2017); Rose and Krich (2021); Do, Gelin, and Tan (2017); Hybl, Albrecht Ferro, and Jonas (2001); Li et al. (2013b); Hedse et al. (2023), suggesting that such factors could have a significant impact in non-perturbative scenarios too. Recently, the ability to isolate specific signals such as the third-order response even in the presence of higher-order contributions was demonstrated in transient absorption experiments Malý et al. (2023). This technique has also been extended to 2DCS experiments performed in the pump-probe geometry Lüttig et al. (2023a, b). Collectively, these observations underscore the need for a comprehensive approach that can accurately simulate nonlinear responses in both perturbative and non-perturbative regimes without relying on simplifying assumptions.
@@
In this work, we present a numerical model based on the phase-cycling method Tian et al. (2003); Tan (2008); Zhang et al. (2012) to calculate the exact nonlinear signal without making assumptions about the pulse envelope or truncating the signal up to a specific order. We illustrate the above capabilities through numerical solutions of the optical Bloch equations (OBEs) for a few-level quantum systems. The calculations are repeated by changing the phase of the excitation pulses and summing them up with specific weights to isolate the desired nonlinear signal. We demonstrate that this approach offers exact solutions for nonlinear signals across both perturbative and nonperturbative regimes without relying on approximations of parameters such as pulse envelopes, pulse ordering, quantum pathways, and order of nonlinear interaction. We begin, in Sec. II, with a generalized description of the phase-cycling method for calculating an arbitrary $N$-wave mixing (NWM) signal, which will be used in the subsequent sections. In Sec. III, we model the saturation of the signal amplitude as the intensity of excitation pulses in 2DCS experiment is increased beyond the $\chi^{(3)}$ limit. We further demonstrate the usefulness of this technique by reproducing previously reported experimental findings in quantum dots through phase cycling simulations in Sec. IV. Finally, we summarize our results in Sec. V along with a discussion of the significance of this work.
@@ II N-wave-Mixing signal
Consider a two-level system (2LS), shown in Fig. 1(a), excited by a sequence of $M$ pulses, which are shown in Fig. 1(b). The $j^{th}$ pulse arrives at time $t=t_{j}$ with time- and position-dependent electric field
$E_{j}(t,\mathbf{r})=G_{j}(t-t_{j})\left[\exp{-i(\omega_{j}t-\mathbf{k}_{j}\cdot\mathbf{r}+\phi_{jM}}+c.c.\right],$ (1)
@@
The light-matter interaction in a 2LS can be described by the optical Bloch equations (OBEs) in the rotating-wave approximation:
@@
${\dot{\rho}}_{01}$ $=-(i\Delta+\gamma_{01})\rho_{01}+\frac{i\mu E_{tot}(t)}{\hbar}(\rho_{00}-\rho_{11}),$ (3b)
where $\rho_{00}$ and $\rho_{11}$ are density matrix elements representing the population in the ground |$\ket{0}$ and excited |$\ket{1}$ states, respectively. $\rho_{01}$ and it’s complex conjugate $\rho^{*}_{01}$ represent the coherence between these states. $\Gamma_{1}$ and $\gamma_{01}$ are population and coherence decay rates, respectively, and detuning $\Delta=(\omega_{01}-\omega_{L})$ where $\omega_{L}$ is the center frequency of excitation pulse. The population terms also satisfy an additional constraint for a closed system: $\rho_{00}+\rho_{11}=1$. Consistent with collinear geometry, we consider all the pulses to be propagating in the same direction; thus, the $\mathbf{k}$, and $\mathbf{r}$ terms have been dropped. The solution to the OBEs gives the exact state of the 2LS after excitation with the pulse sequence. For meaningful interpretation of NWM experiment, a specific signal needs to be isolated.
A perturbative NWM signal can be seen as arising from $|\alpha_{j}|$-order interaction with the $j^{th}$ pulse leading to a total of $N$-order interaction with the pulse sequence. Thus,
@@
and the NWM signal amplitude
$S_{NWM}\propto\prod_{j=1}^{M}|E_{j}|^{|\alpha_{j}|}.$ (5)
The constants $\alpha_{j}$ can be either positive or negative and a collection of these define a specific $N$-wave mixing signal. For example, a phase-matching condition such as
$\phi_{sig}=\sum_{j=1}^{M}\alpha_{j}\phi_{jM}$ (6)
@@
This phase matching condition is equivalent to signal emitted in the direction $\mathbf{k}_{sig}=\sum_{j=1}^{M-1}\alpha_{j}\mathbf{k}_{j}$ for noncollinear excitation schemes.
@@
In the above discussion we have assumed that the phase-cycling steps for each excitation pulse $P_{j}$ is known. However, the choice of the optimal phase-cycling scheme with minimum $N_{PC}$ is not obvious and needs to be determined. Such a phase-cycling scheme must ensure that aliasing of other signal does not happen Tan (2008). We outline a generalized algorithm to perform the above check in Appendix A. This aliasing check can be performed as the values of $P_{j}$ are varied to obtain the optimal phase-cycling scheme computationally.
@@
Finally, we note that the procedure to calculate the NWM signal that we have discussed in this section did not make any approximations regarding the order of interaction, quantum pathways and pulse overlap. Thus, we posit that the phase-cycling procedure can be used to calculate the coherent nonlinear response of the system even in the nonperturbative regime, as long as the equations describing the light-matter interactions are valid. In this respect, this procedure should directly mimic experimental condition, where the nonlinear signal in a particular direction or following a particular phase-matching condition can be detected even when the experiments are performed with intense excitation pulses. As a consequence we can explain results from high-intensity experiments by simply increasing the intensity of excitation pulses. We prove this claim in the following sections by reproducing results from various experiments performed on semiconductor quantum dots (QDs) with excitation beyond the $\chi^{(3)}$ regime. We clarify that a particular NWM signal is technically defined only for the $(N-1)^{th}$ order perturbative nonlinear signal. However, in the rest of this paper we refer to the nonlinear signal satisfying a specific phase-matching condition as the NWM signal even for high-intensity excitation when the perturbative treatment of light-matter interaction is not valid; this convention is consistent with previous work Fras et al. (2016).

### v2

# Coherent Nonlinear Optical Response for High-Intensity Excitation

@@ Abstract
Calculation of the coherent nonlinear response of a system is essential to correctly interpret results from advanced techniques such as two-dimensional coherent spectroscopy (2DCS). Usually, even for the simplest systems, such calculations are either performed for low-intensity excitations where perturbative methods are valid and/or by assuming a simplified pulse envelope, such as a $\delta$-function in time. Here, we use the phase-cycling method for exact calculation of the nonlinear response without making the aforementioned approximations even for high-intensity excitation. We compare the simulation results to several experimental observations to prove the validity of these calculations. The saturation of the photon-echo signal from excitons in a semiconductor quantum well sample is measured. The excitation-intensity dependent measurement shows nonlinear contributions up to twelfth order. Intensity-dependent simulations reproduce this effect without explicitly considering higher-order interactions. Additionally, we present simulation results that replicate previously-reported experiments with high-intensity excitation of semiconductor quantum dots. By accurately reproducing a variety of phenomena such as higher-order contributions, switching of coherent signal, and changes in photon-echo transients, we prove the efficacy of the phase-cycling method to calculate the coherent nonlinear signal for high-intensity excitation. This method would be particularly useful for systems with multiple, well-separated peaks and/or large inhomogeneity.
@@ I Introduction
Nonlinear coherent spectroscopy is a widely used technique for probing complex light-matter interactions at the quantum level Mukamel (1995); Cundiff (2008). A fundamental component of nonlinear spectroscopy is the wave-mixing process Shen (2003); Bokor et al. (1981), where a sequence of laser pulses interacts with the material to generate nonlinear signals based on the relevant nonlinear susceptibility of the material. Among these interactions, four-wave mixing (FWM) is one of the most common, arising from the third-order susceptibility $\chi^{(3)}$ Vaughan et al. (2007). Higher-order susceptibilities $\chi^{(n)}$ lead to wave-mixing processes such as six-wave mixing (SWM) and beyond Zhang et al. (2013); Gibson et al. (1991); Trebino and Rahn (1987); Tominaga and Yoshihara (1996); Steffen and Duppen (1996); Lucht, Trebino, and Rahn (1992); Kang, Hernandez, and Zhu (2004), which have been used to probe the electronic Oliver and Fleming (2015) and vibrational states Bhattacharyya and Fleming (2019); Liu et al. (2019).
@@
Building on the wave-mixing principle, multidimensional coherent spectroscopy (MDCS) extends this approach by measuring the wave-mixing signal field while varying two or more time delays between interacting excitation pulses Li (2023); Oliver (2018); Hamm (2011); Gallagher Faeder and Jonas (1999); Cho (2008); Brixner et al. (2017); Zhang et al. (2023, 2013); Zhang, Wells, and Tan (2012). By performing Fourier transforms along the axes of these time delays, MDCS disentangles the nonlinear interactions of the material across multiple frequency dimensions. This multidimensional approach offers several advantages: it measures spectrally-resolved homogeneous response, isolates individual resonances in congested spectrum, and reveals the dynamic evolution trajectories of states that would otherwise remain hidden in one-dimensional measurements. As a result, MDCS provides in-depth insights into phenomena such as energy transfer Thyrhaug et al. (2016); Bixner et al. (2012); Bressan et al. (2019), chemical reactions Nuernberger, Ruetzel, and Brixner (2015), and quantum coherence processes Butkus et al. (2017); Turner and Nelson (2010).
@@
The spectra obtained from MDCS are rich in information, revealing intricate features like coherent couplings Singh et al. (2014); Li et al. (2013a); Wen and Nelson (2013) and many-body interactions Shacklette and Cundiff (2003); Erementchouk, Leuenberger, and Sham (2007). Proper interpretation of these complex signals requires robust simulations. Conventionally, theoretical and numerical models for simulation rely on a perturbative framework, where the nonlinear response scales linearly with the electric field of each excitation pulse, for weak excitation. In this framework, Double-sided Feynman diagrams (DSFDs) are often employed to map out specific quantum pathways Yang et al. (2007); Rose and Krich (2021); Tekavec, Lott, and Marcus (2007), while response functions integrated with electric fields are used to compute the nonlinear response Mukamel (1995); Gelin et al. (2013). However, the overall response is ultimately truncated to a specific order, as it assumes the perturbative expansion remains valid Gelin, Egorova, and Domcke (2009); Gelin et al. (2003); Süß et al. (2019). Consequently, MDCS experiments are generally conducted using low intensity laser pulses, ensuring that higher-order contributions remain below the noise level of measurements, keeping the analysis within the bounds of perturbation theory.
@@
On the other hand, high-intensity excitation can significantly influence coherent excitation and population transfer dynamics, as higher-order NWM processes become significant. MDCS experiments have demonstrated that intense laser field reveals various interesting phenomena across different systems, such as correlations among multiple particles Liang and Li (2021), coherent control schemes through strong-field coupling Suzuki et al. (2016); Wen and Nelson (2013); Cerullo et al. (1996); Schneider et al. (2011), coherence between highly excited states in quantum wellsTurner and Nelson (2010), and exciton-exciton annihilation (EEA) processes in excitonic systems Malý and Mančal (2018); Trinkunas et al. (2001). Most theoretical descriptions of these phenomena rely on simplified approximations such as approximate electric-field envelopes of the excitation pulses Chen et al. (2017), truncation of the signal up to a particular nonlinear order Hahn and Stock (1998); Bressan and van Thor (2021); Binz et al. (2020); Anda and Cole (2021) and/or neglecting pulse overlap effects Rao, Gelin, and Domcke (2017); Gelin, Egorova, and Domcke (2011), which may not be valid for all cases. Moreover, several studies have highlighted that finite pulse durations and overlaps can affect the MDCS spectra within the perturbative regime Smallwood, Autry, and Cundiff (2017); Rose and Krich (2021); Do, Gelin, and Tan (2017); Hybl, Albrecht Ferro, and Jonas (2001); Li et al. (2013b); Hedse et al. (2023), suggesting that such factors could have a significant impact in non-perturbative scenarios too. Recently, the ability to isolate specific signals such as the third-order response even in the presence of higher-order contributions was demonstrated in transient absorption experiments Malý et al. (2023). This technique has also been extended to 2DCS experiments performed in the pump-probe geometry Lüttig et al. (2023a, b). Collectively, these observations underscore the need for a comprehensive approach that can accurately simulate nonlinear responses in both perturbative and non-perturbative regimes without relying on simplifying assumptions.
@@
In this work, we numerically calculate the exact nonlinear signal for high-intensity excitation using the phase-cycling method Tian et al. (2003); Tan (2008); Zhang et al. (2012). In this approach, we do not make assumptions about the pulse envelope or truncate the signal up to a specific order. We illustrate the above capabilities through numerical solutions of the optical Bloch equations (OBEs) for a few-level quantum systems. The calculations are repeated by changing the phase of the excitation pulses and summing them up with specific weights to isolate the desired nonlinear signal. We demonstrate that this approach offers exact solutions for nonlinear signals across both perturbative and nonperturbative regimes without relying on approximations of parameters such as pulse envelopes, pulse ordering, quantum pathways, and order of nonlinear interaction. We begin, in Sec. II, with a generalized description of the phase-cycling method for calculating an arbitrary $N$-wave mixing (NWM) signal, which will be used in the subsequent sections. In Sec. III, we model the saturation of the signal amplitude as the intensity of excitation pulses in 2DCS experiment is increased beyond the $\chi^{(3)}$ limit. We further demonstrate the usefulness of this technique by reproducing previously reported experimental findings in quantum dots through phase cycling simulations in Sec. IV. Finally, we summarize our results in Sec. V along with a discussion of the significance of this work. We note that methods similar to phase-cycling have been previously used to perform nonperturbative calculations of nonlinear optical response Gelin, Chen, and Domcke (2022), which, to the best of our knowledge, were limited to pump-probe signals Gelin et al. (2013); Gelin, Egorova, and Domcke (2011); Wang and Thoss (2008); Seidner, Stock, and Domcke (1995).
@@ II N-wave-Mixing signal
The signal in a NWM experiment is detected through either the $(N-1)^{th}$-order polarization or the the $N^{th}$-order population Nardin et al. (2015). For example, in a typical FWM experiment, the signal is radiated by the third-order polarization generated by three excitation pulses interacting with the sample. This radiated field is interfered with a local oscillator to detect the phase-resolved signal. Alternatively, a fourth excitation pulse can convert the third-order coherence to a fourth-order population and the desired signal is detected through frequency-domain filtering Tekavec, Lott, and Marcus (2007) or phase-cycling Aeschlimann et al. (2011). The signal obtained for the two detection schemes may not be identical – the population-detected signal may not provide phase information Gelin, Chen, and Domcke (2022), there might be quantitative differences due to ensemble averaging Anda and Cole (2021), and qualitative differences may appear due to the presence of dark states Smallwood et al. (2021). However, these differences are not relevant to the current study; thus, we consider the two detection schemes to be equivalent.
Phase-cycling methods and other simulation approaches Keusters, Tan, and Warren (1999); Gelin, Chen, and Domcke (2022); Seidner, Stock, and Domcke (1995) have previously been formulated for a general class of Hamiltonians. However, for the purposes of this work, we have developed our formalism specifically for a two-level system, as it is most relevant to the systems studied here. This formalism is extended later in the study to few-level systems. We consider a model two-level system (2LS), in Fig. 1(a), which is excited by a sequence of $M$ pulses, shown in Fig. 1(b). The $j^{th}$ pulse arrives at time $t=t_{j}$ with time- and position-dependent electric field
$E_{j}(t,\mathbf{r})=\frac{1}{2}G_{j}(t-t_{j})\left[\exp{-i(\omega_{j}t-\mathbf{k}_{j}\cdot\mathbf{r}+\phi_{jM}}+c.c.\right],$ (1)
@@
The light-matter interaction in a 2LS can be described by the optical Bloch equations (OBEs):
@@
${\dot{\rho}}_{01}$ $=-(i{\omega_{01}}+\gamma_{01})\rho_{01}+\frac{i\mu E_{tot}(t)}{\hbar}(\rho_{00}-\rho_{11}),$ (3b)
where $\rho_{00}$ and $\rho_{11}$ are density matrix elements representing the population in the ground |$\ket{0}$ and excited |$\ket{1}$ states, respectively. $\rho_{01}$ and it’s complex conjugate $\rho^{*}_{01}$ represent the coherence between these states. $\omega_{01}$ and $\mu$ are the transition frequency and dipole moment, respectively. $\Gamma_{1}$ and $\gamma_{01}$ are population and coherence decay rates, respectively. The population terms also satisfy an additional constraint for a closed system: $\rho_{00}+\rho_{11}=1$. Consistent with collinear geometry, we consider all the pulses to be propagating in the same direction; thus, the $\mathbf{k}$, and $\mathbf{r}$ terms have been dropped. We numerically solve the OBEs under the rotating-wave approximation (RWA) Boyd (2008) to obtain the exact state of the 2LS after excitation with the pulse sequence; additional details are provided in the Appendix A. However, for meaningful interpretation of NWM experiment, a specific signal needs to be isolated.
A perturbative NWM signal can be seen as arising from the interaction of a sequence of optical pulses, where each pulse contributes to the signal through a specific interaction order. The amplitude of the population-detected NWM signal
$S_{NWM}\propto\prod_{j=1}^{M}|E_{j}|^{|\alpha_{j}|},$ (4)
where $|E_{j}|$ represents the electric field amplitude of the $j^{th}$ pulse, and $|\alpha_{j}|$ indicates the order of interaction with the $j^{th}$ pulse. The total number of interactions leading to the NWM signal is given by:
@@
The constants $\alpha_{j}$ can be either positive or negative and a collection of these define a specific $N$-wave mixing signal. For example, a phase-matching condition within the framework of the RWA Berman (2011); Mukamel (1995)
$\phi_{sig}=\sum_{j=1}^{M}\alpha_{j}\phi_{jM},$ (6)
@@
Thus, we can rewrite the constraint in Eq. (7) as $\sum_{j=1}^{M-1}\alpha_{j}=1$. This alternate framing explicitly indicates that after $M-1$ pulses, the signal corresponds to a polarization-detected signal of $(N-1)^{th}$ order, and signifies the equivalence of the population-detected signal and the heterodyne-detected radiated-polarization signal Gelin, Chen, and Domcke (2022). Consequently, the phase matching condition defined in Eq. (6) is equivalent to signal emitted in the direction $\mathbf{k}_{sig}=\sum_{j=1}^{M-1}\alpha_{j}\mathbf{k}_{j}$ for noncollinear excitation schemes.
@@
In the above discussion we have assumed that the phase-cycling steps for each excitation pulse $P_{j}$ is known. However, the choice of the optimal phase-cycling scheme with minimum $N_{PC}$ is not obvious and needs to be determined. Such a phase-cycling scheme must ensure that aliasing of other signal does not happen Tan (2008). We outline a generalized algorithm to perform the above check in Appendix B. This aliasing check can be performed as the values of $P_{j}$ are varied to obtain the optimal phase-cycling scheme computationally.
@@
Finally, we note that the procedure to calculate the NWM signal that we have discussed in this section did not make any approximations regarding the order of interaction, quantum pathways and pulse overlap. Thus, we posit that the phase-cycling procedure can be used to calculate the coherent nonlinear response of the system even in the nonperturbative regime, as long as the equations describing the light-matter interactions and the RWA are valid. In this respect, this procedure should directly mimic experimental condition, where the nonlinear signal in a particular direction or following a particular phase-matching condition can be detected even when the experiments are performed with intense excitation pulses. As a consequence, it should be possible to explain results from high-intensity experiments by simply increasing the intensity of excitation pulses in simulations. We prove this claim in the following sections by reproducing results from various experiments performed on semiconductor nanostructures with excitation beyond the $\chi^{(3)}$ regime. We clarify that a particular NWM signal is technically defined only for the $(N-1)^{th}$ order perturbative nonlinear signal. As the excitation-intensity is increased, the NWM signal may not be truly isolated due to the higher-order contributions. For instance, higher-order interactions such as $5^{th}$ and $7^{th}$-order may also contribute to the nonlinear response detected while satisfying the phase-matching condition for a $3^{rd}$-order signal. However, for the sake of brevity, in the rest of this paper we refer to the nonlinear signal satisfying a specific phase-matching condition as the NWM signal even for high-intensity excitation when the perturbative treatment of light-matter interaction is not valid; this convention is consistent with previous work Fras et al. (2016).